OpenStudy (anonymous):

One card is selected from a deck of cards. Find the probability of selecting a black card or a card less than 8. I did it this way: P(B or <8) = P(B) + P(<8) - P(B and <8) = (13/52) + (28/52) - (7/52) = 34/52 = 17/26 The homework key says it's 10/13? What am I doing wrong?

OpenStudy (retireed):

Is an ACE the highest card or the lowest card?

OpenStudy (anonymous):

lowest

OpenStudy (retireed):

Thanks I will start working on this with you but I'm not the best with probabilities.

OpenStudy (welshfella):

P(black) = 1/2 P( <8) = 28/52 = 7/13

OpenStudy (welshfella):

Probability isnt my forte either

OpenStudy (anonymous):

OK, duh! I was thinking of spades. So, that gives me P(B or <8) = P(B) + P(<8) - P(B and <8) = (26/52) + (28/52) - (7/52) = 47/52 Which is still different than the key.

OpenStudy (retireed):

I got 7/13 also. but I'm not sure. I am fairly confident it isn't 10/13.

OpenStudy (anonymous):

Oh crap, I was counting spades again. He's the solution. P(B or <8) = P(B) + P(<8) - P(B and <8) = (26/52) + (28/52) - (14/52) = 40/52 = 10/13

OpenStudy (welshfella):

how did you get 14/52?

OpenStudy (anonymous):

There are 7 spades less than 8 (A, 1,2,3,4,5, 6,7) and 7 clubs. That's 14 cards less than 8.

OpenStudy (welshfella):

AH I see!

OpenStudy (anonymous):

weshfella, your 1/2 above helped me figure out I was miscounting. Thanks!

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

for some reason I was thinking = 8 rather than < 8 !!!!